### v1

# Identifying Spurious Correlations using Counterfactual Alignment

Abstract
Models driven by spurious correlations often yield poor generalization performance. We propose the counterfactual alignment method to detect and explore spurious correlations of black box classifiers. Counterfactual images generated with respect to one classifier can be input into other classifiers to see if they also induce changes in the outputs of these classifiers. The relationship between these responses can be quantified and used to identify specific instances where a spurious correlation exists as well as compute aggregate statistics over a dataset. Our work demonstrates the ability to detect spurious correlations in face attribute classifiers. This is validated by observing intuitive trends in a face attribute classifier as well as fabricating spurious correlations and detecting their presence, both visually and quantitatively. Further, utilizing the CF alignment method, we demonstrate that we can rectify spurious correlations identified in classifiers.
00footnotetext: 1Work done prior to joining Amazon
1 Introduction
Challenges related to neural network generalization and fairness often arise due to covariate shift [19] and shortcut learning [10]. Shortcut learning can lead to models making decisions based on factors not aligned with expectations of the model creator. These powerful models (vision models in this work) leverage a wide array of features and relationships, which may inadvertently incorporate undesired, spurious correlations. These spurious relationships may stem from sample bias (e.g., predicting cows when cows are observed on grass but not on a beach) or may be inherent to the class definition (e.g., predicting cows when an animal has four legs) [6].
Our primary goal is to ensure that models make predictions without relying on spurious correlates, a concept coined as “right for the right reasons” [25]. This means that the features employed for prediction align with expectations of the model creator, rather than features which are spuriously correlated with the target label. Models driven by such spurious correlations often yield poor generalization performance [6, 10]. Consequently, it is crucial to develop methods for uncovering and correcting these unwanted correlations.
Our objective is to understand black-box classifiers, such as pre-trained models, as this is a common use case encountered by practitioners. In this analysis, we utilize counterfactual (CF) images, synthetic images simulating a change in the class label of an image [35]. These synthetic images have features modified such that the prediction of the classifier changes. We can then view the synthetic images to understand the reasons that a prediction was made.
Specifically, we are interested in CF images that are directly generated using the gradients of a classifier [8, 12, 3]. Generating CF images with respect to a classifier is rooted in similar logic to that of crafting adversarial examples. However, the key distinction lies in the constraint that CF images remain within the data manifold of plausible images. The latent space of an autoencoder provides such data manifold. This approach enables us to study the specific features used by a given classifier on a particular input. By keeping the classifier and autoencoder independent, different classifiers can be analysed using a singular fixed autoencoder as a reference point.
The task of interpreting these CF images presents new challenges of scale. Spurious correlations may only exist in a handful of samples where rare features occur together. Locating these samples requires investigating the features used for each prediction, which can be automated using the approaches we present in this study.
The approach we take is to generate CF images for the positive predictions of a classifier over an entire dataset and study the resulting CF images using a collection of different classifiers, which we refer to as downstream classifiers, to observe what outputs are impacted. If the CF images generated for one classifier also change the predictions of other classifiers, we can conclude there is shared feature usage. This shared feature usage can prompt us to investigate unexpected relationships. The relative change metric (Eq. 3) can be used to identify spurious correlations which may occur for a specific example or to identify trends over an entire dataset. We study face attribute classification, recognizing its advantage in allowing visual verification by readers.
Figure 1: Overview of the alignment methodology. An image is encoded, reconstructed, and then processed by a classifier. The counterfactual is generated by subtracting the gradient of the classifier output w.r.t. the latent representation. The resulting representation is reconstructed back into an image. The reconstructed images are processed with multiple classifiers and the classifier outputs can be plotted side by side to study their alignment. The base model value can be used as the x-axis to more easily compare it to the predictions of another classifier. The output changes can be quantified and compared using relative change.
Overall, given the problem setting of spurious correlations, the contributions of this work include:
  • We propose CF alignment to reason about the feature relationships between classifiers. This approach allows for both aggregate quantification and targeted querying of models to identify instances where incorrect reasoning leads to a prediction.
  • Our work demonstrates the ability to detect spurious correlations in face attribute classifiers. This is validated by observing intuitive trends in a face attribute classifier as well as inducing spurious correlations and then detecting their presence, both visually and quantitatively.
  • Utilizing the CF alignment method, we demonstrate that we can rectify spurious correlations identified in classifiers. Experimental results show a bias reduction in 10 out of 12 classifiers.
2 Related Work
Counterfactual (CF) generation can be done in a variety of ways [35]. Classifier specific approaches generally perturb a latent space representation guided by a classifier. Some methods use the gradient of a classifier to guide movement in the latent space by computing the gradient directly [8] or defining a loss that is optimized [12, 3]. Another classifier based approach is to train a model that predicts, using training examples, where to walk in the latent space to change the classifier’s prediction [28]. This has the benefit of working with classifiers that are not differentiable. Another approach is conditional generation which doesn’t use a classifier and instead generates images based on a conditioning variable. This conditioning can be a label [18, 5, 27, 28, 11, 4, 31] or text [7] or by manually adjusting the latent representation [29]. Conditional generation learns a representation from the data and does not capture the exact features utilized by a classifier, as the gradient based methods do.
Several methods exist for studying neural network based classifiers to understand the features they use. Early approaches include gradient based attribute maps (aka saliency heatmaps) that could be overlaid on the image [30]. This was extended to capture more caveats of neural network reasoning[33, 32] but these methods still only generate heatmap explanations. Other work focuses on leveraging occlusions to identify discrete image regions [23]. Other work inspects individual neural network neurons to identify where class information was propagating using linear probing [1] and manual identification [21, 20]. Additional methods focus on concepts represented by layer activations and identifying images that share these representations [14]. Another approach [2] generates synthetic images using human annotators such that the images have known feature similarity (e.g. a face with all features held constant except for skin color). These images are then used to identify spurious correlations between attributes.
3 Alignment Methodology
This work proposes the CF alignment approach as described in Fig. 1. For this approach, we generate counterfactuals using Latent Shift [8]. This approach requires an encoder/decoder model $D(E(x))$ where $E$ is the encoder and $D$ is the decoder, and a classifier $f$ which predicts a target $y$ as follows: $y=f(x)$. The autoencoder and the classifier are trained independently without specific requirements except for being differentiable and operating on the same data domain. To compute a counterfactual, an input image $x$ is encoded using $E(x)$, to produce a latent representation $z$. Perturbations of the latent space are computed for a base classifier $f_{b}$ in Eq. 1 which is then used to produce $\lambda$-shifted samples as shown in Eq. 2. $\lambda$ is determined using an iterative search where the value is decreased in steps until the classifier’s prediction is reduced by 0.6 or begins to increase. $z_{\lambda}=z-\lambda\frac{\partial f_{b}(D(z))}{\partial z}$ (1) $x^{\prime}_{\lambda}=D(z_{\lambda})$ (2)
These counterfactual samples can then be input into downstream classifiers that predict various attributes (e.g. $f_{1}$, $f_{2}$) and observe how their predictions change. Relative change (Eq. 3) allows us to quantify the relationship between a base and downstream classifier, similar to correlation, but accounts for the magnitude of the change. In our experiments, we found correlation would produce false positives when there was only a slight change in the prediction of $f_{1}$ compared to the base classifier $f_{b}$.
$\frac{f_{1}(D(z_{\lambda}))-f_{1}(D(z_{0}))}{f_{b}(D(z_{\lambda}))-f_{b}(D(z_{0}))}$ (3)
4 Experiments
The experiments in this work are performed on the CelebA dataset [17] that contains over 200k celebrity images with 40 facial attribute labels per image. The resolution of the images is 1024$\times$1024.
The pre-trained face classifiers used in this work are sourced from [34]. They were trained on the CelebA dataset [17] to predict 40 different facial attributes on images of dimensions 224$\times$224. We chose this model because it has good performance, is publicly available, and was implemented in PyTorch.
We leverage the VQ-GAN autoencoder from [9] trained on the FacesHQ dataset, which combines the CelebA dataset [17] and the Flickr-Faces-HQ (FFHQ) dataset [13]. The resolution of this model is 256x256.
The Captum library [15] is utilized for baseline attribution methods. PyTorch [22] is used for efficient tensor computation and automatic differentiation. The source code and model weights for all experiments will be released publicly online [^1] https://github.com/ieee8023/latentshift.
Figure 2: Relationships between face attribute classifiers as measured by CF-induced relative changes (left), classifier predictions (middle) and training data labels (right). In (a), base classifiers are along the rows and downstream classifiers are along the columns. Comparing (a) to (b) and (c), shows that many relationships reflected in the CF outputs are preserved from correlations in the training data. The non-symmetric relationship between male and wearing_lipstick, highlighted in green, reflects expected interactions between the two face attribute classifiers. The relationship between male and big_nose, highlighted in red, is strong in both the classifier predictions and ground truth labels but low in CF alignment, indicating that although correlated, these features aren’t exploited by the classifier.
(a) CF alignment for male showing alignment with classifiers related to hair and inverse alignment with classifiers related to makeup. The CF image visually appears to add makeup and remove hair.
(b) CF alignment for attractive showing alignment with classifiers related to makeup, face shape, and age and an inverse alignment with nose size. The CF image visually appears to remove makeup and add bags under the eyes.
(c) CF alignment for heavy_makeup showing alignment with classifiers related to makeup and face shape and inverse alignment with classifiers for nose size and bags under the eyes. The CF image visually shows makeup being removed.
Figure 3: Examples of alignment for specific samples showing the visual counterfactual changes as well as other classifiers with aligned and inverse aligned relationships.
4.1 Aggregate statistics over a dataset
Viewing aggregate CF alignment statistics over a dataset can be useful when investigating a model for bias or spurious correlations. The average relative change between pairs of classifiers for 400 samples per class is shown in Fig. 2a. Figures 2b and 2c show correlations between classifier predictions and the ground truth training labels. Only an interesting subset of classifiers are shown in this figure for clarity, with the full CF alignment matrix in Appendix Fig. 9.
We draw the readers attention to the example relationship between male and big_nose in red, which is high in both classifier prediction and ground truth labels but very low using CF alignment. This indicates that while the attributes (labels) are correlated, the male classifier does not largely use big_nose as a feature to predict male - an expected insight that is verified. A similar insight is found for the class of mouth_slightly_open, which has a relatively strong relationship with smiling.
We also observe many non-symmetric relationships such as male and wearing_lipstick in green, as well as young and attractive. These are also intuitive relationships that we would expect the classifiers to learn. The male classifier likely uses lipstick as a feature to reduce the likelihood that the image contains a male subject (high inverse correlation in Fig. 2c), while the wearing_lipstick classifier would be more focused on the color of the lips. Overall, these results help qualitatively depict that counterfactual reasoning can provide model introspection in a manner that is consistent with human reasoning.
4.2 Studying specific examples
Using the aggregate analysis from §4.1 as a guide, classifiers can be selected to investigate undesired relationships. We will first select a sample which contains an inverse relationship between a base classifier for male and downstream classifier for heavy_makeup in Fig. 2(a). Visually, we can see makeup and lipstick in the CF image and the predictions increase as expected in the classifiers for heavy_makeup, wearing_lipstick, and no_beard. Similarly, there is a decrease in the classifier for goatee. These appear reasonable and provide insight about the features used for prediction. Exploring the more subjective base classifier label attractive in Fig. 2(b) we observe big_nose and bags_under_eyes increase, which can depict human preferences and biases in determining an attractive face.
Figure 4: Example of saliency maps failing to provide meaningful localization when the concepts overlap with each other. Here, saliency map methods localize around the nose which doesn’t provide the ability to distinguish between a big or pointy nose prediction.
4.3 Comparison with saliency maps
Saliency maps for feature attribution could potentially be used to perform a similar analysis by looking at relationships between their generated heatmaps. Here, we present an example that demonstrates their limitation. The attribution methods input gradients, integrated gradients [33], and guided backprop [32] are used to explain the classifier predictions of pointy_nose and big_nose. In Fig. 4 the salient areas for pointy_nose and big_nose are the same and cannot be disambiguated. Comparing these heatmaps would yield a high similarity but we observe in Fig. 2 that these two classifiers have an inverse relationship. Using CF alignment, we can gain a deeper understanding of what features the model is using and better reason about why the decision was made.
4.4 Validation by inducing spurious correlations
In order to further verify the CF alignment approach, we construct a classifier with a known spurious correlation and then demonstrate that this bias is observable in the CF alignment plot. A spurious correlation can be induced in the classifier by composing classifiers:
$f_{biased}(x)=f_{smiling}(x)+0.3f_{arched\_eyebrows}(x)$ (4)
The CF alignment plot for the base smiling classifier predictions in Fig. 4(a) show that arched_eyebrows does not change and we can confirm this in the CF image. The resulting biased classifier can be observed using the arched_eyebrows features in Fig. 4(b) both visually as well as in the CF alignment plot. The CF alignment plot shows that the prediction of arched_eyebrows now changes and is aligned with smiling.
(a) CF for smiling showing eyebrows are unchanged. The horizontal line indicates the prediction of arched_eyebrows is not influenced by the features used for smiling in this image.
(b) CF for the modified smiling classifier with an arched eyebrow spurious correlation showing that eyebrows are now changed.
Figure 5: Example of detecting a spurious correlation in a biased classifier. The classifier is biased with arched eyebrows and this is observed in the alignment plot as well as in the counterfactual image. In order to decrease the prediction of the (now biased) classifier, the CF image can remove the arched eyebrows as well.
4.5 Rectifying bias by composing classifiers
This section demonstrates the use of CF alignment to fix model bias in classifiers. We can reuse the relative change between model predictions as a loss function that can be minimized.
Correcting model bias and spurious correlations is an active area of research. Approaches include averaging model weights [36] or methods of invariant optimization [26, 16, 24, 37]. Composing classifiers can change their response to specific samples that contain relevant features as shown in §4.4. By using this additive composition approach we avoid the variance of model training which makes it easy and reproducible to perform experiments.
First, we demonstrate a single example of correcting bias. In Fig. 6, the CF image and CF alignment plot for the heavy_makeup base classifier show unexpected reductions in lip size. Composing the model with the big_lips classifier using a negative coefficient causes the classifier to lower any feature change focused on the lips as this would increase the prediction of the composed classifier (where the goal of the CF is to decrease it). We can also visually observe other features that change, such as skin color, bushy eyebrows, and the amount that the mouth is open.
(a) As makeup is removed from the face, the lip size also shrinks.
(b) The lip size remains the same as the makeup is removed.
Figure 6: An example of a big_lips spurious correlation being corrected for the heavy_makeup classifier.
Next, this unbiasing is scaled up to a larger number of samples and classifiers. A collection of classifiers that contain varying spurious relationships to a big_nose base classifier (shown in Fig. 7a) are modified to remove their bias until their relative change is as close to 0 as possible. This experiment is performed with a train, validation, and hold out test set in order to demonstrate the generalization of this unbiasing process to unseen data.
The classifiers are modified using a single coefficient ($\beta$) such that the resulting classifier is
$f^{\prime}_{target}(x)=f_{target}(x)+\beta f_{big\_nose}(x)$ (5)
The optimization to compute $\beta$ is performed using a pseudo gradient descent where the gradient is approximated by the mean relative change between the target classifier and the big_nose classifier. By subtracting the relative change from $\beta$, scaled by a learning rate, the relative change ($\psi$) with that classifier will be reduced. All together our training objective (including momentum) is
$\beta_{n}=0.001\psi(f_{target},f_{big\_nose})+0.1\beta_{n-1}$ (6)
We find that using small minibatches of 10 samples works well because the computation time for each CF can take over 1 second on a GPU. Additionally, training with samples which induce a small change in the base classifier during the CF generation process can be challenging. As the classifiers are modified, this base change is reduced, which causes the relative change to become more erratic and prevents the optimization from converging. To prevent this, we use samples with a base change $>0.6$.
The resulting training metrics in Fig. 7 show model biases being summarized and minimized similar to a differentiable training loss. We observe a bias generalization gap between the train, valid, and test sets, indicating some degree of overfitting. Early stopping (red vertical line) is used on the validation set to determine the optimal parameters. We use momentum during training to average over the noise. The resulting reduction in bias is shown in Fig 7b. All biases are considerably reduced except for attractive and big_nose, which could be the result of overlapping features used by both classifiers that cannot be changed.
Figure 7: Training curves showing mean relative change during the optimization of a classifier with respect to the big_nose classifier during each iteration. The red vertical line is the early stopping point based on the validation set. This demonstrates how bias can be quantified and corrected using CF alignment.
Figure 8: CF alignment matrices before and after optimization to remove the bias for big_nose on a hold out test set. The figure shows a dramatic reduction in relative change for most classifiers and big_nose with limited residual impact on other relationships.
5 Limitations
There are limitations and challenges to CF generation [35] where bias can also exist in the CF generation method. Furthermore, although we are optimistic that our method will generalize to additional domains, we focus our current analysis on face attribute classification. Although we do not demonstrate the generalizability of our method to additional domains, using face attribute classification is a deliberate choice enabling qualitative evaluation, in addition to our quantitative evaluation. Additionally, since CF generation relies on a separate autoencoder, improving the representational capability of the autoencoder may improve the fidelity of the CF generation.
6 Conclusion
In this work we propose the counterfactual (CF) alignment approach to reason about the feature relationships in classifiers in aggregate statistics, to locate specific examples with spurious correlations, and propose strategies to mitigate the effect of spurious correlations. These claims are supported with an analysis of face attribute classifiers that identify many expected and unexpected spurious correlations such as attractive with heavy_makeup and heavy_makeup with big_nose. The validity of the CF alignment method is confirmed by inducing and quantifying spurious correlations via additive composition. We then explore rectifying classifiers based on observed spurious correlations. Overall, the proposed approach may serve as an end-to-end or human-in-the-loop system to automatically detect, quantify, and correct spurious correlations for image classification tasks that lead to biased classifier outputs.
7 Acknowledgments
We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results.
Supplementary Material
8 Appendix
Figure 9: The complete matrix of relative changes for all classifiers. Ordering is determined by clustering to group similar classifiers.
(a) pointy_nose
(b) smiling
(c) attractive
Figure 10: Visualization of the training from §4.5 reducing the bias of big_nose. The model being evaluated is $f^{\prime}_{target}(x)=f_{target}(x)+\beta f_{big\_nose}(x)$ where $\beta$ is specified by the coefficient column. The metric here is relative change computed between the $f^{\prime}_{target}(x)$, where $target$ is specified in the caption, and the classifier specified by the column header. The top row is from the training set performed on batches (which explains the variance) and the second row is the results on the entire validation set. The row of each table is for a specific iteration. The iterations proceed downward. The training of attractive shown here does not result in a debiased model, the relative change diverges from 0 indicating a failed training.